OpenStudy (anonymous):

(-4*6) +8/-2 answer(20) *check what i did wrong the choices are: ~ -16 ~ -8 ~8 ~16

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

can u show ur work

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so i did (-4*6) which is -24, then i did -24/-2 which was 12+=20

OpenStudy (mathmath333):

\[(-4*6) +(8/-2 )\\=-24-4\\=-28\]

OpenStudy (mathmath333):

u need to follow the \(PEMDAS\) rule and their must be typo in question or options

Directrix (directrix):

Is @mcnabkt saying that the answer is supposed to be 20?

OpenStudy (mathmath333):

one choice may be \[[(-4*6) +8]/-2\\=(-24+8)/-2=-16/-2=8\]

OpenStudy (mathmath333):

yes

Directrix (directrix):

As it is posted, the answer to the question is the one you posted, @mathmath333

OpenStudy (mathmath333):

i know thats why i said their must be typo

OpenStudy (mathmath333):

as it doesnt ,match the choices

Directrix (directrix):

I think the grouping symbols were not posted correctly. So be it unless the Asker returns.

OpenStudy (anonymous):

i used pemdas its just combing like terms and all

OpenStudy (anonymous):

go to a site called cymath

OpenStudy (anonymous):

alright

OpenStudy (freckles):

you can type exactly what your expression looks like

OpenStudy (freckles):

or draw

OpenStudy (freckles):

and draw any missing ( ) or [ ]

OpenStudy (anonymous):

Yes you can do all that

OpenStudy (freckles):

(-4*6)+8/-2 is not the same as [(-4*6)+8]/-2 by the way

OpenStudy (anonymous):

just try that site

OpenStudy (mathmath333):

yes u can do that ;)

OpenStudy (freckles):

\[\frac{(-4 \cdot 6) +8}{-2} \text{ is not the same as what you posted } (-4 \cdot 6)+\frac{8}{-2}\] but I actually I think @mathmath333 did both expressions for you above

OpenStudy (freckles):

so i won't say anything else :p looks like everything was done

OpenStudy (anonymous):

im using what @juando said

OpenStudy (mathmath333):

can u simplify \(2-4/2-8*1\)

OpenStudy (mathmath333):

ok u need the \(pemdas\) here

OpenStudy (mathmath333):

the multilication and division will occur first \(2-(4/2)-(8*2)\)